A skier starting from rest coasts down a mountain that makes an angle of 25 degrees with the horizontal. The coefficient of kinetic friction between her skis and the snow is 0.200. She coasts for a distance of 10.4 m before coming to the edge of a cliff. Without slowing down, she skis off the cliff and lands dwonhill at a point whose vertical distance is 3.50 m below the edge.

How fast is she going before she lands?

Please draw necessary diagrams and show equations when solving. Thank you!
Answer #1

Using conservation of energy.

Net change in Kinetic energy = Net change in Potential energy + work done by friction.

So 0.5 * m* v^2 = m *g * ( 3.5 + 10.4 * sin 25 ) - 0.2 * m * g* 10.4

So v =sqrt ( 2 * 9.8 * ( 3.5 + 10.4 * sin 25 - 0.2* 10.4 ))

So v = 10.67 m/s